This is tricky for me, A parabola has a vertical axis and passes through the points (0,3), (1,7) and (-2,-17). Find the equation of the parabola. use y=ax^2+by+c
Answer by scott8148(6628)

there are three unknowns; a, b, and c ___ you can generate three equations using the points and y=ax^2+bx+c

(0,3) ___ 3=a(0)^2+b(0)+c ___ 3=c

(1,7) ___ 7=a(1)^2+b(1)+3 ___ 4=a+b ___ 4-a=b

(-2,-17) ___ -17=a(-2)^2+b(-2)+3 ___ -20=4a-2b ___ -20=4a-2(4-a) ___ -20=4a-8+2a ___ -12=6a ___ -2=a ___ 6=b